Solve for the net force experienced by q1. (Show your solution)

 

 

     q1           0.3 m           q2                          0.25  m                q3

          . ___________________. _________________________.

        20 µC                            - 15 µC                                        - 10 µC

 


1
Expert's answer
2021-02-15T17:39:50-0500

According to the Coloumb's law, the force experienced by "q_1" from "q_2" is:


"F_{12} = k\dfrac{|q_1||q_2|}{r_{12}^2}"

where "k = 9\times 10^9 \space N\cdot m^2/C^2" is the Coloumb's constant, "r_{12} = 0.3m" is the distance between "q_1 = 20\times 10^{-6}C" and "q_2 = -15\times10^{-6}C". Thus, obtain:


"F_{12} = 9\times 10^9\cdot \dfrac{20\times 10^{-6}\cdot 15\times 10^{-6}}{0.3^2} =30N"

Similarly, the the force experienced by "q_1" from "q_3" is:


"F_{12} = k\dfrac{|q_1||q_3|}{r_{13}^2}"

where "r_{13} = 0.3m+0.25m = 0.55m" is the distance between "q_1" and "q_3 = -10\times10^{-6}C". Thus, obtain:


"F_{13} = 9\times 10^9\cdot \dfrac{20\times 10^{-6}\cdot 10\times 10^{-6}}{0.55^2} \approx 5.95N"

Since "q_2" and "q_3" have the same sign, both forces are directed in the same direction. Thus, according to the superposition principle, the net force will be:


"F_1 = F_{12} + F_{13} = 35.95N"

Answer. 35.95 N.